# Theoretical Modeling of LLM Self-Improvement Training Dynamics Through Solver-Verifier Gap

### v3

@@ 4 Experiment
This section conducts experiments on self-improvement to verify the theoretical framework. We first introduce experimental setups in Section 4.1. We then present experiment results of the self-improvement process in Section 4.2. We finally explore the differences between solver capability and verifier capability in Section 4.3, showing that the solver-verifier gap generally happens in practice.
@@
4.1 Setup
This section introduces the models, datasets, and key parameters. We consider the following two verification methods: TrueFalse (TF) and Quality Evaluation (QE). We utilize two model families: (a) Phi models: From the Phi family (Abdin et al., 2024), we use Phi-4-mini, Phi-3.5-mini, and Phi-3-mini; (b) Llama models: We use Llama-3.2-3B (Grattafiori et al., 2024) and Llama-3.1-8B. Our experiments mainly focus on the models’ mathematical problem-solving capabilities. Accordingly, we employ two representative datasets: GSM8k (Cobbe et al., 2021) and Math (Hendrycks et al., 2021). In addition, we also consider ProntoQA dataset with TF method in order to evaluate the efficiency of our framework on QA problems. Experimental details are provided in Appendix C.
@@ 4.2 Dynamics of Self-Improvement
This section sketches the dynamics of self-improvement using AI-generated feedback for supervised fine-tuning (SFT), aiming to verify the theoretical framework proposed in Section 3.2. We test the Phi-4-mini, Phi-3.5-mini, Phi-3-mini, and Llama-3.2-3B models. We apply Low-Rank Adaptation (LoRA) (Hu et al., 2022), which significantly reduces the number of updatable parameters, thereby enhancing SFT efficiency. The chosen hyperparameters are detailed in Section C.1.
Results. We present the results of self-improvement on the Phi-4-mini model using the QE method, as depicted in Figure 3. The empirical evidence indicates a consistent enhancement in the accuracy of both the solver and the verifier during the self-improvement process, coupled with a concurrent reduction in their respective uncertainties. Furthermore, a narrowing of the gap $G(t)$ between the solver and verifier is evident. Results for other models and methods are presented in Section C.2.
Validation. To validate our theoretical framework, we fit an exponential model to the uncertainty from $10$ self-improvement epochs. Figure 4 presents the results for the Phi-4-mini model in four different settings (Math / GSM8K datasets train split with QE / TF metrics), which illustrates the evolution of three key metrics: solver uncertainty, verifier uncertainty, and the uncertainty gap. In Figure 4, the exponential model demonstrates a strong fit to the empirical data, with the coefficients of determination ($R^{2}$) exceeding $0.9$. This empirical evidence validates the exponential law proposed in our theoretical work. Results for other models are presented in Section C.2.
@@ 4.3 Verifier Outperforms Solver
In this section, we evaluate the solver and verifier performance of LLM, aiming to validate the utility of the solver-verifier gap used in Section 3. Specifically, the experimental results verify that verifier capability outperforms solver capability consistently. Our evaluation is divided into two settings: self-evaluation and cross-evaluation. Self-evaluation employs the same model for both the solver and verifier roles, whereas cross-evaluation utilizes different models for the solver and verifier.
Self Evaluation. We compare the accuracy and uncertainty of solver and verifier on different models and datasets across different $N$. Given that the verifier selects one response from $N$ candidates generated by the solver, the verifier’s performance is expected to improve as $N$ increases. Detailed results are presented in Section C.2.
Cross Evaluation. To better understand the relationship between solver and verifier capabilities, we also perform cross-evaluation, where one model acts as the solver and a different model acts as the verifier. Results are presented in Section C.2.
@@ 5 Discussions on Cross-Improvement
In this section, we discuss how to model cross-improvement within the framework described in Section 3.2. The key insight is that external data affects the theoretical framework through the verification capability. We first present the theoretical framework in Section 5.1, under the framework with limited external data. We then conduct experiments in Section 5.2 to validate the theoretical findings.
@@ 5.1 Theoretical Framework of Cross-Improvement
In this section, we present the theoretical framework of training dynamics of cross-improvement. We follow the notations in Section 3. Besides, assume that we have limited external data with size $M$. For example, we may acquire $M$ external data in total from a better LLM using API queries.
@@
The parameter $\gamma$ represents the general effect of cross-improvement, while $\eta_{t}$ represents the ratio of external data used in epoch $t$. We assume that the effect $\gamma$ is time invariant, $\gamma>0$ when the verifier that adds external data performs better than the verifier without external data; otherwise $\gamma<0$. Overall, in each step, the cross-data first boosts the verification capability, then it evolves following the mathematical framework in Section 3.2. We illustrate this procedure in Figure 6.
@@
Equation 14 represents the initial conditions, which are the same as Equation 8Equation 15 represents the effects of cross-improvement, where the solver capability after cross-improvement $U_{s}^{c}(t)$ remains unchanged, while the verifier capability increases as discussed in Equation 13. The current capability gap $G^{c}(t)$ is then defined as the gap between the solver capability and the current verifier capability. Equation 17 represents the iterative conditions of cross-improvement. Based on the above formulation, we derive an approximate solution of the ultimate uncertainty:
Proposition 5.1.
@@
Under proposition 5.1, we come to the following conclusions: (i) For cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$, the approximate solution is around the same; (ii) The cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$ outperforms self-improvement with $\sum_{t=1}^{T}\eta_{t}=0$ in terms of solver capability, when $\gamma>0$. The proof of proposition 5.1 is shown in Section B.4.
@@ 5.2 Experiments
In this section, we perform cross-improvement experiments using different allocation strategies.  Figure 6 illustrates the process of cross-improvement. For these experiments, external data are generated by DeepSeek-V3 which achieves accuracy of $60.47\%$ and $91.18\%$ on Math and GSM8k dataset respectively. Specifically, we utilize DeepSeek-V3 responses for fine-tuning instead of BoN responses. We test three allocation strategies: (a) Early: All external data are introduced in the first epoch. (b) Uniform: An equal amount of new external data is introduced in each epoch. (c) Late: All external data are introduced in the eighth epoch.
We present the results on the training data in Table 1. We observe that the average solver accuracy for all three strategies is higher than the baseline on the MATH dataset. However, Phi-4-mini shows a slight performance improvement on GSM8k, while Llama-3.2-3B exhibits a drop on GSM8k dataset. This observation indicates that a key factor influencing the effectiveness of cross-improvement is indeed whether the external model’s verifier capability significantly surpasses that of the original model. In addition, we find that the difference in accuracy between the three strategies is slight, which validates the conclusions under proposition 5.1.
@@ 6 Conclusion
This part provides the supplementary materials. Appendix A provides necessary details omitted before. Appendix B provides omitted derivations of the theoretical framework. Appendix C provides experiment details and omitted experiment results.
@@ Appendix A Supplementary Details
In this section, we provide the details omitted before. Section A.1 presents the omitted illustration. We discuss the difference between our work and prior works, as well as the challenges for self-improvement in Section A.2.
@@ A.1 Omitted illustration
Figure 5 gives an overview of our theoretical framework while Figure 6 illustrates cross-improvement process.
Figure 5: An overview of our theoretical framework.
@@
A.2 Review and Prospect
@@
Different definition: In previous work [Song et al., 2025], the solver capability is defined as the accuracy, while the verifier capability is defined as the accuracy of responses that the model deems good responses. The calculation of accuracy depends on the golden answer contained in the original dataset, which is the most significant difference from our definition. In practice, many datasets do not contain gold answers, so our definition could be more widely applied.
@@ Appendix B Theoretical Derivation
In this section, we present the detailed derivation of proposition 3.1. We first derive Equation 11. From Equation 9 and linear approximation $f(G)\approx kG-b$, we have
@@
The general solution of Equation 19 is
@@
Then we derive Equation 12. From Equation 23, we have
@@
In this section, we present the detailed derivation of corollary 3.1. From proposition 3.1, the final solver uncertainty is given by $U_{s,\infty}=U_{s,0}-\alpha^{\prime}$, where $\alpha^{\prime}=\frac{\alpha}{\alpha-\beta}\delta$ and $\delta=G_{0}-b/k$. Substituting these into the expression for $U_{s,\infty}$ yields $U_{s,\infty}=U_{s,0}-\frac{\alpha}{\alpha-\beta}(G_{0}-\frac{b}{k})$. Assuming $U_{s,0}$ is constant, we have $\frac{\partial U_{s,0}}{\partial G_{0}}=0$. Thus, we have $\frac{\partial U_{s,\infty}}{\partial G_{0}}=-\frac{\alpha}{\alpha-\beta}$.
@@
In this section, we present the detailed derivation of corollary 3.2. The objective of corollary 3.2 is to find the time $t$ such that the capability gap $G(t)$ is within a tolerance $\epsilon$ of its convergence value $G_{\infty}$, which is expressed as:
@@
Starting from the expression for $G(t)$ given in proposition 3.1, we have:
@@
In this section, we present the theoretical derivation of proposition 5.1. According to Section 5.1, the dynamics conditions are formulated as:
@@
With these notations, we can derive the following iteration from Equation 38, 39 and 40:
@@
where $\mathbf{U}(0)$ denotes $\begin{bmatrix}U_{v,0},&U_{s,0},&1\end{bmatrix}^{\top}$. The meaning of $U_{v,0},U_{s,0}$ is defined in Equation 40. Under the circumstances, the following approximation holds true:
@@
The third approximation is based on the approximation $\frac{1}{1+\gamma\eta_{t}}\approx 1-\gamma\eta_{t}$, given $\eta_{t}$ a small quantity. Equation 45 and 46 indicates that we can derive a solution of $U_{s}(T)$ only related to $\sum_{t=1}^{T}\eta_{t}$. Therefore, we come to the conclusion that (i) solver uncertainty at the final epoch $T$ is approximately the same for cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$, and (ii) with the two elements in matrix $\mathbf{\Delta^{\prime}}$ that $\gamma$ appears in both negatively correlated to the term $\gamma\sum_{t=1}^{T}\eta_{t}$, the final solver uncertainty is also negatively correlated to this term. This indicates that cross-improvement with $\gamma>0,\sum_{t=1}^{T}\eta_{t}=1$ outperforms self-improvement with $\gamma>0,\sum_{t=1}^{T}\eta_{t}=0$, in terms of solver capability.
Appendix C Experimental Details
In this section, we detail our experiment hyperparameters and present all results of our experiment. In Section C.1, we detail the hyperparameters chosen in our experiments. We provide all experiment results in Section C.2. All of our experiments are run on 80G NVIDIA A800 GPUs.
C.1 Hyperparameters
In this section, we detail the hyperparameters in Table 2.

### v4

@@ 4 Experiment
This section conducts experiments on self-improvement to verify the theoretical framework. We first introduce experimental setups in Section˜4.1. We then present experiment results of the self-improvement process in Section˜4.2. We finally explore the differences between solver capability and verifier capability in Section˜4.3, showing that the solver-verifier gap generally happens in practice.
4.1 Setup
This section introduces the models, datasets, and key parameters. We consider the following two verification methods: TrueFalse (TF) and Quality Evaluation (QE). We utilize two model families: (a) Phi models: From the Phi family (Abdin et al., 2024), we use Phi-4-mini, Phi-3.5-mini, and Phi-3-mini; (b) Llama models: We use Llama-3.2-3B (Grattafiori et al., 2024) and Llama-3.1-8B. Our experiments mainly focus on the models’ mathematical problem-solving capabilities. Accordingly, we employ two representative datasets: GSM8k (Cobbe et al., 2021) and Math (Hendrycks et al., 2021). In addition, we also consider ProntoQA and MBPP datasets with TF method in order to evaluate our framework on QA and code generation problems. Experimental details are provided in Appendix˜D.
@@ 4.2 Dynamics of Self-Improvement
This section sketches the dynamics of self-improvement using AI-generated feedback for supervised fine-tuning (SFT), aiming to verify the theoretical framework proposed in Section˜3.2. We test the Phi-4-mini, Phi-3.5-mini, Phi-3-mini, and Llama-3.2-3B models. We apply Low-Rank Adaptation (LoRA) (Hu et al., 2022), which significantly reduces the number of updatable parameters, thereby enhancing SFT efficiency. The chosen hyperparameters are detailed in Section˜D.1.
Results. We present the results of self-improvement on the Phi-4-mini model using the QE method, as depicted in Figure˜3. The empirical evidence indicates a consistent enhancement in the accuracy of both the solver and the verifier during the self-improvement process, coupled with a concurrent reduction in their respective uncertainties. Furthermore, a narrowing of the gap $G(t)$ between the solver and verifier is evident. Results for other models and methods are presented in Section˜D.2.
Validation. To validate our theoretical framework, we fit an exponential model to the uncertainty from $10$ self-improvement epochs. Figure˜4 presents the results for the Phi-4-mini model in four different settings (Math / GSM8K datasets train split with QE / TF metrics), which illustrates the evolution of three key metrics: solver uncertainty, verifier uncertainty, and the uncertainty gap. In Figure˜4, the exponential model demonstrates a strong fit to the empirical data, with the coefficients of determination ($R^{2}$) exceeding $0.9$. This empirical evidence validates the exponential law proposed in our theoretical work. Results for other models are presented in Section˜D.2.
@@ 4.3 Verifier Outperforms Solver
In this section, we evaluate the solver and verifier performance of LLM, aiming to validate the utility of the solver-verifier gap used in Section˜3. Specifically, the experimental results verify that verifier capability outperforms solver capability consistently. Our evaluation is divided into two settings: self-evaluation and cross-evaluation. Self-evaluation employs the same model for both the solver and verifier roles, whereas cross-evaluation utilizes different models for the solver and verifier.
Self Evaluation. We compare the accuracy and uncertainty of solver and verifier on different models and datasets across different $N$. Given that the verifier selects one response from $N$ candidates generated by the solver, the verifier’s performance is expected to improve as $N$ increases. Detailed results are presented in Section˜D.2.
Cross Evaluation. To better understand the relationship between solver and verifier capabilities, we also perform cross-evaluation, where one model acts as the solver and a different model acts as the verifier. Results are presented in Section˜D.2.
@@
4.4 Impact of Training Stages on Self Improvement
In this section, we investigate how upstream training stages influence the self-improvement dynamics, we conducted a controlled comparative analysis. This section addresses the potential confounder that models at different stages might exhibit distinct optimization behaviors. We utilize the EvoLM model suite Qi et al. (2025) for this study, which allows us to control for model architecture (Llama-2 1B) while isolating the impact of training stages. We compared three checkpoints representing the full lifecycle:
  • Standard SFT Baseline (Base Stage): Derived from the raw pre-trained model (1B-160BT) via standard SFT. This represents a model with high latent potential but no domain-specific mid-training.
  • Mid-train Enhanced Model (Mid-train Stage): Derived from a model that underwent domain-specific Continued Pre-training (CPT) on math data (1B-160BT-8+42BT) before SFT. This represents the domain-primed state.
  • Fully Post-trained Model (Post-train Stage): A model that has completed the full pipeline, including mid-training, SFT and Reinforcement Learning (1B-160BT-8+42BT-100Kep1-100Kep16). This represents the converged state.
We present the results in Figure˜5 and Figure˜9. The results shows a clear evolution in training dynamics governed by model training stage: Base model exhibit the highest plasticity, driven by the largest initial gap. Mid-training lowers the initial potential energy and increases initial capability. Post-trained models demonstrate dynamic saturation, validating the asymptotic convergence as the capability gap is minimized.
@@ 5 Discussions on Cross-Improvement
In this section, we discuss how to model cross-improvement within the framework described in Section˜3.2. The key insight is that external data affects the theoretical framework through the verification capability. We first present the theoretical framework in Section˜5.1, under the framework with limited external data. We then conduct experiments in Section˜5.2 to validate the theoretical findings.
Figure 5: Self-improvement dynamics across Base, Mid-trained, and Post-trained stages on train split. Fitted curves ($R^{2}$) reveal a clear evolutionary trajectory: (1) The Base-SFT model exhibits the highest Solver Rate ($\alpha=146.3$) and Decay Rate ($0.244$). (2) The Mid-trained model reduces the initial uncertainty, proving that mid-training optimizes the initialization state. (3) The Post-trained model shows the lowest rate ($\alpha=16.0$), confirming the theory of saturation near the capability limit.
@@ 5.1 Theoretical Framework of Cross-Improvement
In this section, we present the theoretical framework of training dynamics of cross-improvement. We follow the notations in Section˜3. Besides, assume that we have limited external data with size $M$. For example, we may acquire $M$ external data in total from a better LLM using API queries.
@@
The parameter $\gamma$ represents the general effect of cross-improvement, while $\eta_{t}$ represents the ratio of external data used in epoch $t$. We assume that the effect $\gamma$ is time invariant, $\gamma>0$ when the verifier that adds external data performs better than the verifier without external data; otherwise $\gamma<0$. Overall, in each step, the cross-data first boosts the verification capability, then it evolves following the mathematical framework in Section˜3.2. We illustrate this procedure in Figure˜7.
@@
Equation˜13 represents the initial conditions, which are the same as Equation˜7Equation˜14 represents the effects of cross-improvement, where the solver capability after cross-improvement $U_{s}^{c}(t)$ remains unchanged, while the verifier capability increases as discussed in Equation˜12. The current capability gap $G^{c}(t)$ is then defined as the gap between the solver capability and the current verifier capability. Equation˜16 represents the iterative conditions of cross-improvement. Based on the above formulation, we derive an approximate solution of the ultimate uncertainty:
Proposition 5.1
@@
Under Proposition˜5.1, we come to the following conclusions: (i) For cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$, the approximate solution is around the same; (ii) The cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$ outperforms self-improvement with $\sum_{t=1}^{T}\eta_{t}=0$ in terms of solver capability, when $\gamma>0$. The proof of Proposition˜5.1 is shown in Section˜B.4.
@@ 5.2 Experiments
In this section, we perform cross-improvement experiments using different allocation strategies.  Figure˜7 illustrates the process of cross-improvement. For these experiments, external data are generated by DeepSeek-V3 which achieves accuracy of $60.47\%$ and $91.18\%$ on Math and GSM8k dataset respectively. Specifically, we utilize DeepSeek-V3 responses for fine-tuning instead of BoN responses. We test three allocation strategies: (a) Early: All external data are introduced in the first epoch. (b) Uniform: An equal amount of new external data is introduced in each epoch. (c) Late: All external data are introduced in the eighth epoch.
We present the results on the training data in Table˜1. We observe that the average solver accuracy for all three strategies is higher than the baseline on the MATH dataset. However, Phi-4-mini shows a slight performance improvement on GSM8k, while Llama-3.2-3B exhibits a drop on GSM8k dataset. This observation indicates that a key factor influencing the effectiveness of cross-improvement is indeed whether the external model’s verifier capability significantly surpasses that of the original model. In addition, we find that the difference in accuracy between the three strategies is slight, which validates the conclusions under Proposition˜5.1.
@@ 6 Conclusion
Ethics Statement
This research is theoretical and methodological, focusing on developing a mathematical model for the training dynamics of LLM self-improvement. All experiments were conducted on public academic benchmarks and did not involve human subjects or sensitive data, thus requiring no Institutional Review Board (IRB) approval. Our work aims to deepen the understanding of the self-improvement mechanism to foster the development of more predictable and reliable models.
Reproducibility Statement
To ensure the reproducibility of our work, we provide detailed descriptions of our theoretical and experiment setup. Our theoretical model is detailed in Section˜3, with complete mathematical proofs for all propositions and corollaries provided in Appendix˜B. The experiment setup is described in Section˜4.1. Further implementation details, such as the self-improvement algorithm (Algorithm˜1) and the hyperparameters used to obtain the results (Table˜3), are available in Appendix˜D.
Acknowledgments
We would like to thank the members of the Jiaye’s Group for their insightful discussions throughout this project. We are also grateful to the anonymous reviewers for their constructive comments which helped improve the paper. This work is supported by Shanghai Science and Technology Development Funds 24YF2711700 and Fundamental Research Funds for the Central Universities 2024110586. Xin Liu is partially supported by the National Natural Science Foundation of China 12201383 and Innovative Research Team of Shanghai University of Finance and Economics. Yifan Sun would like to express his deepest gratitude to his beloved girlfriend, Jinjin Zheng, for her unwavering companionship and support, which served as a constant source of strength throughout the inevitable lows of this research journey. Yushan Liang would like to extend her special thanks to her parents for their unwavering support.
@@ Acknowledgments
This part provides the supplementary materials. Appendix˜A provides necessary details omitted before. Appendix˜B provides omitted derivations of the theoretical framework. We claim the use of LLMs in Appendix˜CAppendix˜D provides experiment details and omitted experiment results.
@@ Appendix A Supplementary Details
In this section, we provide the details omitted before. Section˜A.1 presents the omitted algorithm and illustration, while Section˜A.2 summarizes main contributions of our work. In Section˜A.3, we discuss the relationship between uncertainty and capability. Additionally, we further discuss the difference between our work and prior works, as well as the challenges for self-improvement in Section˜A.4.
@@ A.1 Omitted illustration
Figure˜6 gives an overview of our theoretical framework. Algorithm˜1 presents one-step self-improvement used in our experiments while Figure˜7 illustrates cross-improvement process.
Figure 6: Overview of the theoretical framework on self-improvement and cross-improvement.
Algorithm 1 Self Improvement (One Step)
[…]
2:for each prompt $x_{i}\in\mathcal{X}$ do
3:  $C_{x_{i}}\leftarrow\emptyset$;
4:  for $j\leftarrow 1$ to NN do
5:   Generate response $\hat{y}_{i,j}\sim\pi_{f}(\cdot|x_{i})$;
6:   Ask model to evaluate $\hat{y}_{i,j}$ and return a score $s(\hat{y}_{i,j})$;
7:   if $s(\hat{y}_{i,j})\geq\sigma$ then
8:    Append $\hat{y}_{i,j}$ to $C_{x_{i}}$;
9:   end if
10:  end for
11:  $\hat{y}_{i}^{\text{BoN}}\leftarrow\operatorname*{arg\,min}_{C_{x_{i}}}\frac{1}{L(\hat{y}_{i,j})}U_{f}(\hat{y}_{i,j}|x_{i})$;
12:  Append $\hat{y}_{i}^{\text{BoN}}$ to $\mathcal{Y}^{\text{BoN}}$;
13:  Append $U_{f}(\hat{y}_{i}^{\text{BoN}})$ to $\mathcal{U}^{\text{BoN}}$;
14:end for
[…]
16: Self-improved model $\hat{f}$
@@
A.2 Contributions
In this section, we summarize main contributions of our work. Our main contributions are:
  • Theoretical framework: We detail our theoretical framework on the dynamics of self-improvement in Section˜3, inspired by the potential energy framework based on the solver-verifier capability gap in Equation˜6. Theoretical derivations in Equation˜11 imply an exponential law for the model capability, and we further establish the extreme capability based on the framework.
  • Experiments: We conduct experiments on self-improvement to verify the theoretical framework in Section˜4 across multiple models, datasets, and verification methods. Our empirical observations indicate that (i) the uncertainty/accuracy during the self-improvement process follows an exponential law with an extreme capability (Figure˜3) as implied by the theoretical framework; (ii) the solver-verifier gap generally happens in practice across different regimes (Section˜4.3), validating the utility of the gap as the potential energy.
  • Cross-improvement: We further investigate in Section˜5 the cross-improvement under the above framework, where limited external data is provided during the training process. We contend in Section˜5.1 that cross-improvement might improve the verifier capability, thus improving the extreme capability of self-improvement. Empirically, we observe in Section˜5.2 that if the verifier capability is improved by the external data, the performances of LLM are enhanced, and vice versa (Table˜1). This accords with the theoretical findings.
A.3 relationship between uncertainty and capability
In this section, we clarify the relationship between uncertainty and capability. Although low uncertainty is often correlated with high capability (i.e., accuracy), they are not conceptually equivalent. In this work, we use uncertainty instead of accuracy as a proxy metric for model capability. The justifications are detailed as follows:
  • Sharpening Mechanism: Following Huang et al. (2025), minimizing the loss ($U_{v}$) forces the model to sharpen its intrinsic probability distribution onto high-quality responses. A decrease in $U_{s}$ signifies that this distributional shift has occurred and that the model has successfully learned from the verifier’s output. Thus, $U_{s}$ is the direct measure of the model’s capability.
  • Empirically correlation between accuracy and uncertainty: To validate that $U_{s}$ indicates the model’s capability, we calculate the Pearson correlation coefficient (r) between $U_{s}$ and solver accuracy for different models and datasets (Table˜2). Accuracy and Uncertainty for most model-dataset pairs are significantly negatively correlated ($|r|>0.8$), which proves that reducing solver uncertainty ($U_{s}$) translates to gains in accuracy.
  • Absence of Ground Truth: In many real-world scenarios, particularly open-ended generation tasks, ground-truth labels are often unavailable or difficult to define. Unlike accuracy, which strictly depends on external gold references, uncertainty is computed from the model’s output distribution. This makes uncertainty a necessary and universal metric for monitoring capability evolution in label-free regimes where accuracy calculation is infeasible.
  • The Optimization Objective: In standard SFT, the objective of training is to maximize the likelihood of correct responses, which is mathematically equivalent to minimizing their uncertainty (Negative Log-Likelihood). In our framework, we regard the output of the verifier as the ground truth for SFT. The loss function at step $t$ is defined as $L_{t}(f)\triangleq U_{v}(t)=-\frac{1}{n}\sum_{i=1}^{n}\log\ \pi_{f}(\hat{y}_{i}^{\text{BoN}}(t)|x_{i})$, which is identical to our definition of verifier uncertainty $U_{v}(t)$.
Table 2: Correlation coefficients between Uncertainty and Accuracy.
Dataset Method Llama-3.2-3B Phi-3-mini Phi-3.5-mini Phi-4-mini
GSM8k QE $0.963^{***}$ $0.985^{***}$ $0.974^{***}$ $0.991^{***}$
TF $66^{***} 0.990$-$90^{***}$ $0.967^{***}$ $0.988^{***}$
Math QE $0.965^{***}$ $0.958^{***}$ $0.177$ $0.983^{***}$
TF $0.967^{***}$ $0.900^{***}$ $0.167$ $0.978^{***}$
ProntoQA TF $0.867^{***}$ $0.897^{***}$ $0.764^{***}$ $0.850^{***}$
MBPP TF $0.865^{***}$ $0.951^{***}$ $0.936^{***}$ $0.6735^{*}$
Note: $^{***}\,p<0.001$, $^{**}\,p<0.01$, $^{*}\,p<0.05$.
A.4 Review and Prospect
@@
Different definition: In previous work (Song et al., 2025), the solver capability is defined as the accuracy, while the verifier capability is defined as the accuracy of responses that the model deems good responses. The calculation of accuracy depends on the golden answer contained in the original dataset, which is the most significant difference from our definition. In practice, many datasets do not contain gold answers, so our definition could be more widely applied.
@@ Appendix B Theoretical Derivation
In this section, we present the detailed derivation of Proposition˜3.1. We first derive Equation˜10. From Equation˜8 and linear approximation $f(G)\approx kG-b$, we have
@@
The general solution of Equation˜18 is
@@
Then we derive Equation˜11. From Equation˜22, we have
@@
In this section, we present the detailed derivation of Corollary˜3.1. From Proposition˜3.1, the final solver uncertainty is given by $U_{s,\infty}=U_{s,0}-\alpha^{\prime}$, where $\alpha^{\prime}=\frac{\alpha}{\alpha-\beta}\delta$ and $\delta=G_{0}-b/k$. Substituting these into the expression for $U_{s,\infty}$ yields $U_{s,\infty}=U_{s,0}-\frac{\alpha}{\alpha-\beta}(G_{0}-\frac{b}{k})$. Assuming $U_{s,0}$ is constant, we have $\frac{\partial U_{s,0}}{\partial G_{0}}=0$. Thus, we have $\frac{\partial U_{s,\infty}}{\partial G_{0}}=-\frac{\alpha}{\alpha-\beta}$.
@@
In this section, we present the detailed derivation of Corollary˜3.2. The objective of Corollary˜3.2 is to find the time $t$ such that the capability gap $G(t)$ is within a tolerance $\epsilon$ of its convergence value $G_{\infty}$, which is expressed as:
@@
Starting from the expression for $G(t)$ given in Proposition˜3.1, we have:
@@
In this section, we present the theoretical derivation of Proposition˜5.1. According to Section˜5.1, the dynamics conditions are formulated as:
@@
With these notations, we can derive the following iteration from Equation˜37, 38 and 39:
@@
where $\mathbf{U}(0)$ denotes $\begin{bmatrix}U_{v,0},&U_{s,0},&1\end{bmatrix}^{\top}$. The meaning of $U_{v,0},U_{s,0}$ is defined in Equation˜39. Under the circumstances, the following approximation holds true:
@@
The third approximation is based on the approximation $\frac{1}{1+\gamma\eta_{t}}\approx 1-\gamma\eta_{t}$, given $\eta_{t}$ a small quantity. Equation˜44 and 45 indicates that one can derive a solution of $U_{s}(T)$, which is only related to $\sum_{t=1}^{T}\eta_{t}$. Therefore, one can come to the conclusion that (i) solver uncertainty at the final epoch $T$ is approximately the same for cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$, and (ii) with the two elements in matrix $\mathbf{\Delta^{\prime}}$ that $\gamma$ appears in both negatively correlated to the term $\gamma\sum_{t=1}^{T}\eta_{t}$, the final solver uncertainty is also negatively correlated to this term. This indicates that cross-improvement with $\gamma>0,\sum_{t=1}^{T}\eta_{t}=1$ outperforms self-improvement with $\gamma>0,\sum_{t=1}^{T}\eta_{t}=0$, in terms of solver capability.
Appendix C The Use of Large Language Models
We utilize Large Language Models (LLMs) to enhance the clarity, conciseness, and grammatical accuracy of the paper. The model’s suggestions help refine sentence structure and improve the overall readability of the paper.
Appendix D Experimental Details
In this section, we detail our experiment hyperparameters and present all results of our experiment. In Section˜D.1, we detail the hyperparameters chosen in our experiments. We provide all experiment results in Section˜D.2. All of our experiments are run on 80G NVIDIA A800 GPUs.
D.1 Hyperparameters
In this section, we detail the hyperparameters in Table˜3.